# GenVC: Self-Supervised Zero-Shot Voice Conversion

### v1

@@ Abstract
Zero-shot voice conversion has recently made substantial progress, but many models still depend on external supervised systems to disentangle speaker identity and linguistic content. Furthermore, current methods often use parallel conversion, where the converted speech inherits the source utterance’s temporal structure, restricting speaker similarity and privacy. To overcome these limitations, we introduce GenVC, [^1] Audio samples are available at https://caizexin.github.io/GenVC/index.html. The code and model checkpoints will be released shortly. a generative zero-shot voice conversion model. GenVC learns to disentangle linguistic content and speaker style in a self-supervised manner, eliminating the need for external models and enabling efficient training on large, unlabeled datasets. Experimental results show that GenVC achieves state-of-the-art speaker similarity while maintaining naturalness competitive with leading approaches. Its autoregressive generation also allows the converted speech to deviate from the source utterance’s temporal structure. This feature makes GenVC highly effective for voice anonymization, as it minimizes the preservation of source prosody and speaker characteristics, enhancing privacy protection.
@@
Voice Conversion, Large Language Model, Speech Anonymization, Speech Synthesis, Speech Generation
@@ I Introduction
Voice conversion (VC) aims to modify the vocal characteristics of a source signal while preserving its linguistic content. This technology has numerous valuable applications, including assisting individuals with speech impairments, enhancing privacy through voice anonymization, and facilitating voice dubbing for films. Progress in VC has closely mirrored advancements in text-to-speech synthesis, with recent methods achieving remarkable naturalness, producing converted voices perceptually indistinguishable from authentic human recordings [1]. Despite these advances, significant challenges remain, particularly in cloning novel voices and adapting to diverse recording conditions. These limitations primarily arise from the difficulty of training scalable and robust models capable of effectively handling such variability. Furthermore, leading VC approaches typically perform parallel conversion of linguistic features into corresponding acoustic features. Although these systems effectively modify acoustic characteristics such as timbre, this parallel conversion process often results in converted utterances that retain the temporal and prosodic structure of the source speaker [2, 3]. This retained similarity can expose characteristics of the source speaker, diminishing the naturalness and potentially compromising the intended voice transformation [4].
Early VC research primarily relied on parallel data, where models were trained using paired source and target speech recordings with identical linguistic content [5]. While this simplifies the VC modeling process, collecting such data is often labor-intensive, costly, and impractical for large-scale recording efforts. These limitations also reduce the generalizability of VC models in challenging yet practical scenarios, such as multi-speaker or cross-lingual VC [6]. In contrast, non-parallel VC approaches eliminate the need for paired utterances, greatly expanding VC’s applicability to a wider range of scenarios, including zero-shot VC, which enables the cloning of an unseen speaker’s voice even if they are not included in the training set [7, 8, 9].
In particular, zero-shot VC requires two components for speech disentanglement: one to capture the latent representation of a speaker’s voice and another to extract linguistic content. To this end, many studies utilize external models, such as automatic speech recognition (ASR), automatic speaker verification (ASV), or text-to-speech (TTS) to facilitate the disentanglement process [10, 11]. These external models require labeled datasets for effective training to be useful for VC. For example, an ASV system should be trained on a speaker-rich corpus to accurately represent distinct voices [12]. This raises a fundamental research question: Can a VC system be trained without any supervision? In addition, previous zero-shot VC methods struggle to generalize to unseen speakers [13, 14]. A promising direction to address this limitation is to scale up data utilization to cover a more diverse range of voice types, thereby improving generalization [15].
In this paper, we propose Generative Voice Conversion (GenVC), a novel approach for zero-shot VC that eliminates the need for external supervised models. Our method achieves state-of-the-art performance in speaker similarity and privacy preservation, while maintaining competitive naturalness compared to leading VC approaches. The key contributions of this work are as follows:
  • To the best of our knowledge, GenVC is the first zero-shot VC method that does not rely on external supervised models like ASR, ASV and TTS.
  • Self-supervised disentanglement of speaker characteristics and linguistic content enables GenVC to be trained on more diverse datasets, leading to enhanced voice cloning performance.
  • An autoregressive modeling approach allows zero-shot VC to overcome a common limitation in previous VC models, where the converted speech was constrained to match the duration of the source utterance.
@@ II Related Work
VC can be regarded as a subset of speech generation tasks. The advent of self-supervised learning (SSL) models and language models (LMs), known for their strong performance in context modeling and scalability in natural language processing, has inspired their application to sequential tasks in the audio domain.
@@
Concurrently, neural audio tokenizers, or speech codecs, have enabled the conversion of audio into discrete units while preserving high-fidelity reconstruction from low-bitrate tokens, making them well-suited for LM-based frameworks [16]. These advancements have driven the development of LM-based speech generation models, such as AudioLM [17]. AudioLM utilizes tokens generated by a SoundStream codec [18] and employs a hierarchical modeling framework involving three LMs to sequentially predict semantic, coarse acoustic, and fine acoustic tokens. This approach allows for the generation of natural and coherent audio continuations from short input prompts. Similarly, AudioGen integrates LMs and audio tokens to generate audio content conditioned on descriptive text inputs [19].
@@
These advancements in LM-based speech generation have significantly influenced the development of generative zero-shot text-to-speech (TTS) models [20, 21, 22]. VALL-E [23, 24], one of the earliest LM-based TTS models, utilizes Encodec [25] as the audio tokenizer and incorporates both auto-regressive and non-autoregressive LMs to predict audio tokens from input text or phonemes. Building upon this foundation, subsequent iterations of VALL-E have explored tasks such as cross-lingual modeling and improved token representation for greater efficiency [26, 27]. Beyond purely LM-based frameworks, certain scalable TTS approaches have integrated additional methodologies. For instance, CosyVoice combines LMs for text-to-token generation with conditional flow-matching models for token-to-speech synthesis, creating a hybrid system [28]. Similarly, MaskGCT extends the LM and tokenization paradigm by replacing the auto-regressive mechanism with non-autoregressive, masking-based generation strategies, offering an alternative approach to efficient audio synthesis [29].
Nevertheless, most codecs utilized by LM-based generation systems produce a fixed number of codes—typically 8—per time frame, complicating prediction due to the parallel generation of multiple codes at each step. While strategies such as delaying codec prediction have been proposed to mitigate this challenge [30, 31], the high number of codes still adds computational complexity and increases the difficulty of modeling. In contrast, scalable approaches like Tortoise-TTS [15] and XTTS [32] overcome these limitations by discretizing audio into a single token per frame. This simplification not only reduces architectural complexity but also enhances compatibility with LM training.
Research on LM-based approaches for zero-shot VC remains limited. Building on AudioLM, Wang et al. introduced LM-VC [33], which employs three LMs to perform semantic and acoustic modeling. This framework supports large-scale training and achieves impressive performance in terms of speaker similarity and speech naturalness. However, like AudioLM, its complex architecture requires all three models during inference, leading to slow processing and rendering it unsuitable for streaming conversion. StreamVoice explored the streaming capabilities of zero-shot LM-based VC to address these limitations [34]. Unlike its predecessors, StreamVoice uses only a single LM, simplifying the architecture. However, it relies on an externally trained, supervised ASR model to guide semantic information inference. Additionally, the generated speech duration directly matches that of the source speech, preserving the prosody structure of the source utterance. While this design maintains prosody consistency, it also introduces source-speaker information leakage in the converted speech, a limitation shared with many traditional VC systems.
@@ III GenVC
Inspired by prior research in speech generation, we present GenVC, a zero-shot generative VC model. Our approach integrates discrete audio tokenizers with a generative Transformer-based LM for VC, simplifying the complexity of earlier LM-based VC methods. This design enables the model to be trained entirely in an self-supervised manner, supporting large-scale training and maximizing the utilization of available data. Furthermore, the autoregressive nature of the model allows for the conversion of source utterances into the voice and style of a target speaker without preserving the prosodic structure of the source utterances. This capability ensures that our model produces natural-sounding utterances comparable to leading zero-shot VC models while also making it suitable for privacy-sensitive applications.
III-A Overview
Figure 1 provides an overview of the proposed generative VC system. At the core of this approach is a causal Transformer-based LM that predicts a sequence of acoustic tokens conditioned on both a fixed-length voice style representation and a sequence of linguistic tokens. Two auxiliary modules, DVAEs and vocoder, manage the compression, discretization, and reconstruction of audio signals. The VC process unfolds in three sequential phases, each building upon the previous one and optimized through reconstruction losses, eliminating the need for labeled data or external supervised models:
@@ III-A Overview
    Generative Modeling: An decoder-only LM that autoregressively generates audio tokens conditioned on acoustic style representations and linguistic tokens.
@@
The subsequent sections provide detailed descriptions of the models and methodologies employed in each phase.
@@
Tokenization is a crucial step in LMs for breaking down inputs into smaller, manageable units. In the context of VC, the objective is to transform the speaker’s timbre (style) of a source utterance while preserving its original linguistic content. To achieve this, tokenization is employed to extract, compress, and normalize features from audio signals for effective model training. In our VC approach, two distinct types of tokens are used: 1. Phonetic Tokens: Capturing the linguistic content of the audio, representing the “what” is being said. 2. Acoustic Tokens: Encoding the acoustic properties of the audio, including timbre, prosody, environmental background, and other stylistic elements.
These tokens are derived from different audio representations. Studies have shown that self-supervised learning (SSL) speech models can consistently and significantly produce representations enriched with phonetic information [35]. In the context of VC, SSL models serve as ideal feature extractors for capturing the phonetic units of a given speech signal [36, 37]. For a given input utterance $x\in\mathbb{R}^{T}$, the phonetic embedding sequence extracted using a pre-trained SSL model is denoted as $\mathbf{O}\in\mathbb{R}^{T_{o}\times d_{o}}$, where $T_{o}$ represents the sequence length, and $d_{o}$ is the feature dimensionality. For acoustic representations, spectrograms are commonly used to capture the complex properties of audio signals. Such feature can be extracted and represented as $\mathbf{A}\in\mathbb{R}^{T_{a}\times d_{a}}$, where $T_{a}$ denotes the sequence length, and $d_{a}$ specifies the dimensionality of the acoustic features.
As shown in Figure 1, our approach utilizes discrete variational autoencoders (DVAEs) [38] to compress $\mathbf{O}$ into phonetic tokens $[\mathbf{o}_{1},\dots,\mathbf{o}_{n}],\mathbf{o}_{i}\in 1,2,\dots,K_{p}$ and $\mathbf{A}$ into acoustic tokens $[\mathbf{a}_{1},\dots,\mathbf{a}_{m}],\mathbf{a}_{i}\in 1,2,\dots,K_{a}$. Here, $n$ and $m$ denote the lengths of the tokenized sequences derived from the input feature sequences, while $K_{o}$ and $K_{a}$ represent the predefined number of discrete codes for the phonetic and acoustic tokenizers, respectively. The DVAE architecture follows those used in Tortoise-TTS [15] and XTTS [32] and is detailed in Appendix A.
@@
The backbone model in Phase 2 is an LM built on a decoder-only Transformer architecture [39]. This phase also incorporates a Perceiver encoder to extract fixed-length style representations from the prompt utterance [40]. As depicted in Figure 1, the causal language model predicts acoustic tokens by conditioning on both the style prompt derived from the Perceiver encoder and the phonetic tokens extracted from the source audio clip.
The Perceiver encoder processes a variable-length sequence of acoustic features extracted from the audio prompt and generates a fixed-dimensional audio style representation. This is achieved by using a set of learned latent vectors as queries, while the keys and values are formed by concatenating these latent vectors with the acoustic features extracted from the prompt audio. The Perceiver encoder architecture consists of cross-attention blocks, detailed in Appendix A. The output of the Perceiver encoder $\mathbf{E}_{\text{style}}\in\mathbb{R}^{T_{s}\times d_{\text{model}}}$ follows the distribution $P(\mathbf{E}_{\text{style}}\mid\mathbf{A}_{\text{prompt}},\mathbf{E}_{\text{latent}};\theta_{\text{Perceiver}})$ where $\mathbf{E}_{\text{latent}}$ is the learnable latents that have the same shape as $\mathbf{E}_{\text{style}}$. $T_{s}$ denotes the fixed number of latent sequence, and $d_{\text{model}}$ is the dimensionality of the LM.
@@
$\prod_{t=1}^{n}P(\mathbf{o}_{i}\mid\mathbf{o}_{s:i-1},\mathbf{E}_{\text{style}};\theta_{\text{LM}})$ (1)
@@
$\prod_{j=1}^{m}P(\mathbf{a}_{j}\mid\mathbf{a}_{s:j-1},\mathbf{o}_{s:e},\mathbf{E}_{\text{style}};\theta_{\text{LM}}).$ (2)
@@
Here, $\alpha$ and $\beta$ denote the weights assigned to the phonetic and acoustic token prediction losses, respectively. The phonetic token loss, $L_{\text{phonetic}}$ is calculated as the negative log-likelihood of the distribution defined in Expression 1, while the acoustic token loss, $L_{\text{acoustic}}$ is the negative log-likelihood of the distribution defined in Expression 2. Since the primary task of the VC system aligns with that of TTS systems—predicting acoustic tokens—we follow the approach outlined in  [32] and assign $\beta$ a significantly larger value than $\alpha$ ($\beta\gg\alpha$).
@@
Phase 3 focuses on reconstructing the audio waveform from the generative predictions. Since Phase 2 outputs an acoustic token sequence where each token $\mathbf{a}_{i}$ is highly compressed due to vector quantization, directly reconstructing audio from these acoustic codes often introduces pronunciation issues and artifacts [32]. To mitigate these issues, we utilize a HiFiGAN [41] vocoder conditioned on features $\mathbf{H}\in\mathbb{R}^{m\times d_{\text{model}}}$ derived from the final hidden layer of the LM. Unlike systems such as XTTS, which rely on additional speaker embeddings obtained from a separate neural module, we find that conditioning the HiFiGAN vocoder solely on the LM’s final hidden features is sufficient for producing high-quality audio reconstructions.
III-B Training and Inference
During training, as shown in Figure 1, both the audio prompt and the audio clip are randomly segmented from the same source utterance. Typically, the linguistic content of the audio prompt differs from that of the audio clip. This setup allows the Perceiver encoder to capture acoustic attributes that are not represented by the linguistic tokens. Additionally, it facilitates the disentanglement of linguistic content from speaker-specific information in the input audio clip. As a result, the encoder effectively learns to extract both acoustic environment characteristics and speaker-specific representations from the audio prompt. Notably, this approach eliminates the need for supervision from an external speaker embedding model, which is a common feature of traditional zero-shot VC methods. Meanwhile, phonetic and acoustic tokens are derived from the audio clip, and the input sequence to the LM is structured as follows:
$\scriptsize[<\text{style embeddings}><\text{phonetic tokens}><\text{acoustic tokens}>]$
During inference, a significant portion of the model can be omitted, including the decoder of the phonetic tokenizer, the entire acoustic tokenizer, and the discriminator. The inference process, given a source utterance $\textbf{Audio}_{\text{src}}$ and a target utterance $\textbf{Audio}_{\text{tgt}}$ specifying the target voice, proceeds as follows:

### v2

@@ Abstract
Most current zero-shot voice conversion methods rely on externally supervised components, particularly speaker encoders, for training. To explore alternatives that eliminate this dependency, this paper introduces GenVC, a novel framework that disentangles speaker identity and linguistic content from speech signals in a self-supervised manner. GenVC leverages speech tokenizers and an autoregressive, Transformer-based language model as its backbone for speech generation. This design supports large-scale training while enhancing both source speaker privacy protection and target speaker cloning fidelity. Experimental results demonstrate that GenVC achieves notably higher speaker similarity, with naturalness on par with leading zero-shot approaches. Moreover, due to its autoregressive formulation, GenVC introduces flexibility in temporal alignment, reducing the preservation of source prosody and speaker-specific traits, and making it highly effective for voice anonymization. [^1] Audio samples, code, and model checkpoints are available at
https://caizexin.github.io/GenVC/index.html
@@
Voice Conversion, Language Model, Speech Anonymization, Speech Synthesis, Speech Generation
publicationid: pubid: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. This work is accepted by the IEEE ASRU 2025.
@@ I Introduction
Zero-shot Voice conversion (VC) seeks to transform a source voice to match an unseen target speaker, with minimal adaptation, while preserving the original linguistic content [1, 2, 3]. Progress in this area has closely followed advancements in zero-shot text-to-speech (TTS) synthesis, with recent models achieving remarkable naturalness—generating speech that is perceptually indistinguishable from that of real human speakers [4]. Despite these gains, significant challenges remain, particularly in cloning novel voices and adapting to diverse recording conditions. These limitations arise primarily from the difficulty of training robust and scalable models capable of handling such variability. Furthermore, leading VC approaches operate by converting linguistic features into acoustic representations in a parallel manner. While these systems effectively modify acoustic traits such as timbre, the parallel conversion process often preserves the source speaker’s temporal and prosodic patterns [5, 6, 7, 8]. As a result, the converted speech retains perceptual cues from the original speaker, which diminishes both the naturalness and the similarity of the intended voice transformation [9, 10].
Achieving high-quality zero-shot VC typically requires disentangling speaker identity and linguistic content through two dedicated modules: one that captures vocal characteristics and the other that extracts linguistic content. To achieve this separation, many existing approaches leverage pre-trained models, such as automatic speech recognition (ASR), automatic speaker verification (ASV), or TTS, which rely on supervised training with labeled datasets [11, 12]. Minimizing the level of supervision in training, however, enhances the scalability of VC systems and allows the exploitation of extensive unlabeled speech corpora [13, 14]. One promising strategy for improving zero-shot synthesis involves scaling data to include a broader range of voice types, thereby enhancing generalization to unseen speakers [15]. Minimally supervised frameworks like NANCY [13] demonstrate the feasibility of this direction, but often rely on frame-aligned mapping, which can inadvertently preserve source utterance’s temporal and prosodic structure. In contrast, autoregressive architectures provide a compelling alternative to model temporal dependencies more faithfully, and produce prosodic patterns that better reflect the target speaker’s style [16].
To address these challenges, we propose GenVC, a generative zero-shot VC system designed with three key objectives: (1) reducing dependency on external supervision through self-supervised learning for disentangling speaker and linguistic features, thereby enhancing scalability; (2) leveraging an autoregressive generation mechanism to better model target speaker style and improve voice similarity; and (3) maintaining controllability by encoding speaker characteristics into a compact high-dimensional space, enabling applications such as voice anonymization. To achieve these goals, GenVC employs speech tokenization and is built upon a causal Transformer-based architecture. Experimental results show that GenVC achieves significant improvements in speaker similarity for unseen VC tasks and enhanced privacy preservation in anonymization evaluations, while maintaining naturalness competitive with leading VC methods.
@@ II Related Work
VC can be viewed as a subset of speech generation tasks. The rise of self-supervised learning (SSL) models and language models (LMs), known for their strengths in contextual modeling and scalability in natural language processing, has prompted their adoption in sequential audio tasks. Concurrently, neural audio tokenizers, also referred to as speech codecs, have enabled the transformation of continuous audio signals into discrete tokens while maintaining high-fidelity reconstruction. These audio tokenizers are particularly well-suited for integration into LM-based architectures [17, 18], paving the way for a new wave of LM-driven speech generation models. One leading example is AudioLM [19], which leverages SoundStream [20] for tokenization and employs a hierarchical arrangement of three LMs to predict semantic, coarse acoustic, and fine acoustic tokens in sequence. This approach allows for the generation of natural and coherent audio continuations from short prompts.
@@
These advancements in speech generation have significantly influenced the development of generative zero-shot TTS systems [21, 22, 23, 24, 25]. Among the earliest examples is VALL-E [26], which employs Encodec [27] as the audio tokenizer and incorporates both auto-regressive and non-autoregressive LMs to predict audio tokens from input text or phonemes. Later iterations of VALL-E extended its capabilities to tasks such as cross-lingual synthesis and more efficient token representation [28, 29]. Beyond purely LM-based models, hybrid systems have also emerged. For example, CosyVoice [30] combines LM-based text-to-token generation with conditional flow-matching models for token-to-speech synthesis. Another model, MaskGCT [31], extends the LM and tokenization paradigm by replacing the auto-regressive mechanism with non-autoregressive, masking-based generation strategies for efficient audio synthesis [31].
A persistent challenge in LM-based speech generation lies in the nature of current speech codecs, which typically produce multiple tokens—often eight—per frame. This parallel token prediction increases modeling complexity and computational cost. Techniques like delayed token prediction [32, 33] have been proposed to address this issue, yet the overhead remains significant. In contrast, models such as Tortoise-TTS [15] and XTTS [34] adopt a simpler approach, discretizing audio into a single token per frame. This design choice not only streamlines the model architecture but also aligns more effectively with standard LM training paradigms.
Research on LM-based approaches for zero-shot VC remains limited. Building on the AudioLM framework, Wang et al. proposed LM-VC [16], which incorporates three LMs to model semantic and acoustic features. While the model demonstrates strong performance in preserving speaker identity and generating natural-sounding speech, its reliance on a multi-component architecture results in slow inference and makes it unsuitable for real-time applications. To address this, StreamVoice [35] introduces a simplified architecture using a single LM to enable streaming zero-shot VC. However, it depends on a separately trained supervised ASR model to extract semantic content. Furthermore, the output speech duration is fixed to match the source speech, preserving prosodic features but also leading to speaker information leakage from the source—an issue shared with many traditional VC systems.
A concurrent LM-based approach, Vevo [36], addresses similar limitations by introducing a controllable framework for timbre and style conversion. The system comprises two stages: an autoregressive transformer followed by a flow-matching transformer. Both stages are trained with self-supervised, in-context learning, making the framework scalable. However, Vevo conditions speaker representation directly on the source spectrogram rather than using a high-dimensional embedding. While this design is effective for voice imitation, it limits the model’s ability to generate pseudo voices, reducing its suitability for applications like speech anonymization [37].
@@ III GenVC
Inspired by prior research in speech generation, we present GenVC, a zero-shot generative VC model. Our approach integrates discrete audio tokenizers with a generative Transformer-based LM for VC, simplifying the complexity of earlier LM-based VC methods. This design enables the model to be trained entirely in an self-supervised manner, supporting large-scale training. Furthermore, the autoregressive nature of the model allows for the conversion of source utterances into the voice and style of a target speaker without preserving the prosodic structure of the source utterances.
Figure 1 presents an overview of GenVC. At its core, GenVC incorporates a causal Transformer-based LM designed to produce acoustic token sequences conditioned on a fixed-length voice style embedding and a sequence of linguistic tokens. Two auxiliary components, the speech tokenizer and vocoder, handle the compression, discretization, and reconstruction of audio. The voice conversion process unfolds in three sequential stages, each building on the last and trained via reconstruction losses, without requiring labeled data or external supervision.
@@
    Generative Modeling: A decoder-only LM that autoregressively generates audio tokens conditioned on acoustic style representations and linguistic tokens.
@@
Tokenization is a crucial step in LMs for breaking down inputs into smaller, manageable units. In the context of VC, the objective is to transform the speaker’s timbre/style of a source utterance while preserving its original linguistic content. To achieve this, two distinct types of tokens are used in GenVC: 1. Phonetic Tokens: Capturing the linguistic content of the audio, representing “what” is being said. 2. Acoustic Tokens: Encoding the acoustic properties of the audio, including timbre, prosody, environmental background, and other stylistic elements.
These tokens are derived from different audio representations. Studies have shown that SSL speech models can consistently and significantly produce representations enriched with phonetic information [38]. In the context of VC, SSL models serve as ideal feature extractors for capturing the phonetic units of a given speech signal [39, 40]. For a given input utterance $x\in\mathbb{R}^{T}$, the phonetic embedding sequence extracted using a pre-trained SSL model is denoted as $\mathbf{O}\in\mathbb{R}^{T_{o}\times d_{o}}$, where $T_{o}$ represents the sequence length, and $d_{o}$ is the feature dimensionality. For acoustic representations, spectrograms are commonly used to capture the complex properties of audio signals. Such features can be extracted and represented as $\mathbf{A}\in\mathbb{R}^{T_{a}\times d_{a}}$, where $T_{a}$ denotes the sequence length, and $d_{a}$ specifies the dimensionality of the acoustic features.
As shown in Figure 1, our approach utilizes discrete variational autoencoders (DVAEs) [41] to compress $\mathbf{O}$ into phonetic tokens $\mathbf{o}_{1},\dots,\mathbf{o}_{n}],\mathbf{o}_{i}\in 1,2,\dots,K_{p}$ and $\mathbf{A}$ into acoustic tokens $\mathbf{a}_{1},\dots,\mathbf{a}_{m}],\mathbf{a}_{i}\in 1,2,\dots,K_{a}$. Here, $n$ and $m$ denote the lengths of the tokenized sequences derived from the input feature sequences, while $K_{o}$ and $K_{a}$ represent the predefined number of discrete codes for the phonetic and acoustic tokenizers, respectively. The DVAE architecture follows those used in Tortoise-TTS [15] and XTTS [34].
@@
The backbone model in Phase 2 is an LM built on a decoder-only transformer architecture [42]. This phase also incorporates a Perceiver encoder to extract fixed-length style representations from the prompt utterance [43]. As depicted in Figure 1, the causal LM predicts acoustic tokens by conditioning on both the style prompt derived from the Perceiver encoder and the phonetic tokens extracted from the source audio clip. Note that the audio clip here refers to a segment of the source utterance.
The Perceiver encoder processes a variable-length sequence of acoustic features extracted from the audio prompt and generates a fixed-dimensional audio style representation [44]. This is achieved by using a set of learned latent vectors as queries, while the keys and values are formed by concatenating these latent vectors with the acoustic features extracted from the prompt audio. The Perceiver encoder architecture consists of cross-attention blocks. The output of the Perceiver encoder $\mathbf{E}_{\text{style}}\in\mathbb{R}^{T_{s}\times d_{\text{model}}}$ follows the distribution $P(\mathbf{E}_{\text{style}}\mid\mathbf{A}_{\text{prompt}},\mathbf{E}_{\text{latent}};\theta_{\text{Perceiver}})$ where $\mathbf{E}_{\text{latent}}$ is the learnable latents that have the same shape as $\mathbf{E}_{\text{style}}$. $T_{s}$ denotes the fixed number of latent sequence, and $d_{\text{model}}$ is the dimensionality of the LM.
@@
$\prod_{t=1}^{n}P(\mathbf{o}_{i}\mid\mathbf{o}_{s:i-1},\mathbf{E}_{\text{style}};\theta_{\text{LLM}})$ (1)
@@
$\prod_{j=1}^{m}P(\mathbf{a}_{j}\mid\mathbf{a}_{s:j-1},\mathbf{o}_{s:e},\mathbf{E}_{\text{style}};\theta_{\text{LLM}}).$ (2)
@@
Here, $\alpha$ and $\beta$ denote the weights assigned to the phonetic and acoustic token prediction losses, respectively. The phonetic token loss, $L_{\text{phonetic}}$ is calculated as the negative log-likelihood of the distribution defined in Expression 1, while the acoustic token loss, $L_{\text{acoustic}}$ is the negative log-likelihood of the distribution defined in Expression 2. Since the primary task of the VC system aligns with that of TTS systems—predicting acoustic tokens—we follow the approach outlined in  [34] and assign $\beta$ a significantly larger value.
@@
Phase 3 focuses on reconstructing the audio waveform from the generative predictions. Since Phase 2 outputs an acoustic token sequence where each token $\mathbf{a}_{i}$ is highly compressed due to vector quantization, directly reconstructing audio from these acoustic codes often introduces pronunciation issues and artifacts [34]. To mitigate these issues, we utilize a HiFiGAN [45] vocoder conditioned on features $\mathbf{H}\in\mathbb{R}^{m\times d_{\text{model}}}$ derived from the final hidden layer of the LM. Unlike systems such as XTTS, which rely on additional speaker embeddings obtained from a separate neural module, we found that conditioning the HiFiGAN vocoder solely on the LM’s final hidden features is sufficient for producing high-quality audio reconstructions.
III-A Training and Inference
During training, as shown in Figure 1, both the audio prompt and the audio clip are extracted from the same source utterance using random start points and segment lengths. Phonetic and acoustic tokens are derived from the audio clip. Typically, the linguistic content of the audio prompt differs from that of the audio clip. This setup allows the Perceiver encoder to capture acoustic attributes that are not represented by the linguistic tokens. Additionally, it facilitates the disentanglement of linguistic content from speaker-specific information in the input audio clip. As a result, the encoder effectively learns to extract both acoustic characteristics and speaker-specific representations from the audio prompt. This approach eliminates the need for supervision from an external speaker embedding model, which is a common feature of traditional zero-shot VC methods.
During inference, auxiliary components can be omitted, such as the phonetic tokenizer decoder, the entire acoustic tokenizer, and the discriminator. The inference process, given a source utterance $\textbf{Audio}_{\text{src}}$ and a target utterance $\textbf{Audio}_{\text{tgt}}$ specifying the target voice, proceeds as follows: